find all the factors of 48 then classify 48 as primee or composite
Answer by jim_thompson5910(35256)

The factors of 48 are:

1, 2, 3, 4, 6, 8, 12, 16, 24, 48.

Now lets find the prime factors:

Solved by pluggable solver: Find factors of any number
48 is NOT a prime number: 48 = 2 * 2 * 2 * 2 * 3

Work Shown

48 is divisible by 2: 48 = 24 * 2.
24 is divisible by 2: 24 = 12 * 2.
12 is divisible by 2: 12 = 6 * 2.
6 is divisible by 2: 6 = 3 * 2.
3 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.



So this means 48 is a composite number